OpenStudy (anonymous):

Solve the proportion x/x +1 = x /x+4. What is the value(s) of x?

OpenStudy (anonymous):

cross multiplie first

OpenStudy (mertsj):

What do you get when you cross multiply?

OpenStudy (anonymous):

x(x+4)=x(x+1)

OpenStudy (cwrw238):

x^2 + 4x = x^2 + x 4x = x! hmm..

OpenStudy (anonymous):